

********************************************************************
UARS WEEKLY STATUS REPORT - 19 APR 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues under day only operations due to low power
conditions. We are performing default solar observations during this
time.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM support of the UARS Traceability Mission for the TIMED mission has
been organized with the GVUI team. Predicted TIMED-UARS conjunctions from
the high geographic latitudes were compared with the expected magnetic position
and the PEM climatology. Positional and anticipated precipitation flux
information showed that the expected electron precipitation as a function of
magnetic activity had the highest probably of a successful conjunction on
April 19. Continued monitoring has shown changes in conjunction times. For
conjunctions within 300 km, there are now 4 times shown:
Contact # Date Day Time Geo Lat ILAT MLT Hemisphere
(deg) (deg) (hr)
1 April 18 108 21:30 56.87 54.43 2.96 North
2 April 18 108 22:18 -56.89 66.06 6.07 South
3 April 23 113 14:56 56.98 53.09 2.77 North
4 April 23 113 15:44 -56.98 48.03 2.24 South
Contact times have remained about the same as from the last update; however,
the time offsets and interspacing of spacecraft have changed a bit which has
caused the last contact on April 18 to no longer meet the requirement above.
Of coarse, what is important is that PEM passes through the aurora at the
same time GUVI is imaging at the same location, and this does not necessarily
mean that UARS and TIMED are in conjunction.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 April 2002 (Orbit 57784, GMT Day 096) through 12
April 2002 (Orbit 57889, GMT Day 102).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 100/1859 57856 Power OFF (YAW)
PEM 101/1858 57871 Power OFF
SUSIM 101/1858 57871 Power OFF
The beta angle ranged between the angles of 32.82 degrees to 7.14
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 13 April 2002 (DOY 103).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.68 to 0.95 degrees C. Additional
battery performance data for each battery on 05 April 2002 and 12
April 2002 are provided in Appendix D.
The clock error ranged from -10.1 to 14.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HALOE = 01
PEM AXIS = 02
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 0
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
All FOT efforts are currently being directed to the readiness of
Eclipse for the upcoming yaw maneuver on 13 April 2002. All converted
CECIL procedures are being re-executed against the simulator for all
instrument turn ons, as well as all of the POCC resident microloads
involved in the turn ons. The yaw script and yaw command sheets have
been modified to reflect Eclipse command syntax. The stored command
load has been checked and the FOT is ready for the yaw.
MEETINGS
The FOT attended a UARS YAW Maneuver CAM on Thursday, April 11, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 13 April 2002 (DOY 103). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 April 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57853 100/1404 02-119 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
57856 100/1859 02-120 HALOE POWER OFF (YAW)
57869 101/1402 02-121 PEM CALIBRATION
57871 101/1858 02-122 PEM POWER OFF
57871 101/1858 02-123 SUSIM POWER OFF
57875 102/0118 02-124 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
57880 102/0956 02-125 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 April 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 April 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57853 100/1407 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
57875 102/0128 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
57880 102/1000 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 April 2002
05 April 2002 - GMT Day 095 Beta = 34.23 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.16/-4.38 +0.0/ +0.0 9.6 0.000 0.0
2 + 11.2/- 16.8 +4.48/+3.82 +26.8/-13.2 26.4 1.040 18.9
3 + 11.2/- 22.4 +3.82/+2.85 +25.6/-12.8 26.4 1.030 18.5
12 April 2002 - GMT Day 102 Beta = 7.14 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.16/-4.38 +0.0/ +0.0 0.0 0.000 0.0
2 + 11.2/+134.0 +4.50/+3.87 +27.8/-15.8 25.0 1.010 20.6
3 + 22.4/- 22.4 +3.55/+2.60 +26.7/-15.8 25.0 1.010 20.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 April 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-023 FOT 57767 094/2027 24285 Early LOS on OMNI Pass
02-024 FOT 57783 095/2209 24285 SOLSTICE INOPS in Day